OpenStudy (anonymous):

How do I get the surface area and volume for a sphere to t nearest whole number?

OpenStudy (campbell_st):

volume of a sphere, with radius r is \[V = \frac{4}{3} \pi r^3\] the surface area is \[SA = 4 \pi r^2\]